OpenStudy (anonymous):

Evaluate the triple integrals: ∫∫∫xydV, where Q is the solid tetrahedron with Q vertices (0,0,0), (1,0,0), (0,2,0), and (0,0,3). Please explain step by step

OpenStudy (anonymous):

Here is an example straight from Calc book. EXAMPLE 2 Finding the Limits of Integration in the Order dy dz dx Set up the limits of integration for evaluating the triple integral of a function F(x, y, z) over the tetrahedron D with vertices (0, 0, 0), (1, 1, 0), (0, 1, 0), and (0, 1, 1). Solution We sketch D along with its “shadow” R in the xz-plane (Figure 15.29). The upper (right-hand) bounding surface of D lies in the plane The lower (left-hand) bounding surface lies in the plane The upper boundary of R is the line The lower boundary is the line First we find the y-limits of integration. The line through a typical point (x, z) in R parallel to the y-axis enters D at and leaves at Next we find the z-limits of integration. The line L through (x, z) parallel to the z-axis enters R at and leaves at Finally we find the x-limits of integration. As L sweeps across R, the value of x varies from to The integral is \[\int\limits_{0}^{1}\int\limits_{0}^{1-x}\int\limits_{x+z}^{1}F(x,y,z) dy dz dx\] =\[\int\limits_{0}^{1}\int\limits_{0}^{1-x}\int\limits_{x+z}^{1} dy dz dx\] =\[\int\limits_{0}^{1}\int\limits_{x}^{1}(y-x) dy dx\] =\[\int\limits_{0}^{1}[(1/2)y^2-xy]_{y=x}^{y=1}dx\] =\[\int\limits_{0}^{1}(1/2-x+(1/2)x^2) dx\] =\[[(1/2)x-(1/2)x^2+1/6x^3]_{0}^{1}\] =1/6 Make sure to make a drawing so that you can see where the lines intersect and where it projects onto x plane. Hope this helps.

OpenStudy (anonymous):

can you draw a graph? i understand how to integrate for the most part i just dont know how to get the limits for the integrals

OpenStudy (anonymous):

The graph is in the attached file.

OpenStudy (anonymous):

how would i get the equation from the graph?

OpenStudy (anonymous):

the tetrahedron equation?

OpenStudy (anonymous):

You need to find the equations of some sides in their planes and the equation of the plane (isoceles triangle)

OpenStudy (anonymous):

What is the equation of the plane (isoceles triangle)?

OpenStudy (anonymous):

i dont know how to get the equation of the plane

OpenStudy (anonymous):

y= x+ z

OpenStudy (anonymous):

how did you get y = x+ for the equation of an isosceles triangle?

OpenStudy (anonymous):

x+z*

OpenStudy (anonymous):

You may have to read about planes and lines and know this subject well. This will help you understand double and triple integration.